OpenStudy (anonymous):

Calculus help please Solve the problem: The sales in thousands of a new type of product are given by s(t)=100-60e^-0.8t when t=3 I would think the s'(3)=144 thousand, but the answer is 4.4 thousand

zepdrix (zepdrix):

\[\large\rm s(t)=100-60e^{-0.8t}\]\[\large\rm s'(t)=0-60e^{-0.8t}\cdot(-0.8)\]

zepdrix (zepdrix):

Derivative of 100 is simply 0. Maybe that's what was making your answer so large. Hmm, forget chain rule also maybe?

OpenStudy (anonymous):

I forget chain rule. Thank you